The base of a triangle sail is 4x feet and its height is 3x + 10 feet. Write a binomial in terms of x for the area of the sail.
Answer by richwmiller(17219)

1/2b*h=a
1/2*(4x)*(3x+10)=a
2x*(3x+10)=a
6x^2+20x=a